OpenStudy (anonymous):

Please help. Write a rule for the nth term of the sequence. Use your rule to find a100.–8, 9, 26, 43, 60, . . .

OpenStudy (anonymous):

Your sentence seems scrambled... "Use your rule to find a100.–8" What's going on there?

OpenStudy (anonymous):

Use your rule to find a100. The sequence is –8, 9, 26, 43, 60, . . .

OpenStudy (anonymous):

OK look at the difference between the consecutive numbers in the sequence. How much between -8 and 9? How much between 9 and 26? Do you see a pattern?

OpenStudy (anonymous):

Yeah, I see an increase of 17

OpenStudy (anonymous):

Good. So your rule "a" will look like: a = 17n + b where n is the "nth" term in the sequence b is some number. To solve for b, take the first number in the sequence (n=1, 1st number). When n=1, a=-8. So: a = 17n + b -8 = 17*1 + b Can you solve for b?

OpenStudy (anonymous):

Umm... I'm not really sure. - 8 = 17*1 + b = - 8 = 17 + b. subtract 17 from both sides? - 25 = b?